# A GLOBAL MINIMUM TAX FOR LARGE FIRMS ONLY:

### v6

@@
A GLOBAL MINIMUM TAX FOR LARGE FIRMS ONLY:
IMPLICATIONS FOR TAX COMPETITION***Andreas Haufler, Seminar for Economic Policy, LMU Munich, Akademiestr. 1, 80799 Munich, Germany. Email: Andreas.Haufler@econ.lmu.de.  Hayato Kato, Graduate School of Economics, the University of Osaka, 1-7 Machikaneyama, Toyonaka, Osaka 560-0043, Japan. Email: hayato.kato@econ.osaka-u.ac.jp.
 Financial support from the German Research Foundation through CRC TRR 190 (Haufler) and the JSPS KEKENHI (JP20H01495, JP22K13388, JP24H00014) (Kato) is gratefully acknowledged. This study is conducted as a part of the Project “Economic Policy Issues in the Global Economy” undertaken at the Research Institute of Economy, Trade and Industry (RIETI). Kato thanks the University of Munich for their hospitality while part of this paper was written. We thank Tomohiro Ara, Real Arai, Ruby Doeleman, Antoine Ferey, Emanuel Hansen, Jim Hines, Niels Johannesen, Makoto Hasegawa, Jota Ishikawa, Mick Keen, Sang-Hyun Kim, Yoshimasa Komoriya, Jonas Löbbing, Mohammed Mardan, Jakob Miethe, Kenichi Nishikata, Hirofumi Okoshi, Pascalis Raimondos, Dirk Schindler, Karl Schulz, Georg Thunecke, and Alfons Weichenrieder for many helpful comments and discussions. Thanks also to seminar and conference participants at the IIPF in Utah State, MPI Berlin, Goettingen, Osaka, Kyoto, Shimane, QUT Brisbane, Ryukoku, Gakushuin, CESifo (Munich & Venice Summer Institute), and RIETI for helpful comments.
Andreas Haufler and Hayato Kato
@@ A GLOBAL MINIMUM TAX FOR LARGE FIRMS ONLY:
The Global Minimum Tax (GMT) is applied only to firms above a certain size threshold, permitting countries to set differential tax rates for small and large firms. We analyse tax competition among multiple tax havens and a non-haven country for heterogeneous multinationals to evaluate the effects of this partial coverage of GMT. Upon the introduction of a moderately low GMT rate, the havens commit to the single uniform GMT rate for all multinationals. However, gradual increases in the GMT rate induce the havens, and subsequently the non-haven, to adopt discriminatory, lower tax rates for small multinationals. Our calibration exercise shows that introducing a GMT rate of 15% results in a regime where only the havens adopt split tax rates. Welfare and tax revenues fall in the havens but rise in the non-haven, yielding a positive net gain worldwide.
@@
Profit shifting by multinational enterprises (MNEs) remains a central challenge for international corporate taxation. Aggregate estimates suggest that more than one-third of foreign-earned corporate profits are shifted to tax havens (Tørsløv et al., , 2023), a pattern also documented in microdata for France (Davies et al., , 2018), the UK (Bilicka, , 2019), and the US (Guvenen et al., , 2022) In response to the resulting revenue losses, the OECD launched its Base Erosion and Profit Shifting (BEPS) initiative (OECD, , 2013), whose core outcome is the introduction of a global minimum tax (GMT), also known as Pillar Two. [^1] See Fuest and Neumeier, (2023); Hebous et al., (2024) for the institutional background. Another outcome of the BEPS initiative is Pillar One, which reallocates taxing rights from the country of a MNE’s residence to the market jurisdiction (Mukunoki et al., , 2025; Ogawa and Tsuchiya, , 2025; Richter, , 2025). In 2021, more than 130 countries including many tax havens agreed on a 15% GMT; by August 2025, 65 countries had drafted or enacted legislation implementing the GMT. [^2] See Tax Foundation, ‘The Latest on the Global Tax Agreement’: https://taxfoundation.org/blog/global-tax-agreement/ (accessed October 1, 2025). Despite announcing after the June 2025 G7 summit that it would not fully comply with the GMT, the US had already implemented a similar tax on Global Intangible Low-Tax Income under the 2017 Tax Cuts and Jobs Act (Chodorow-Reich et al., , 2024).
@@
A limitation of the GMT is that it applies only to ‘large’ multinationals, defined as groups with at least 750 million EUR in annual revenues in two of the last four years. [^3] Since 2016, such groups have been required to file country-by-country reports on their activities, profits, and taxes in each country of operation (Doeleman et al., , 2024, Section 2). The revenue threshold reflects the balance between limiting profit shifting and minimising compliance costs (OECD, , 2022, p.14). According to the Orbis database of Bureau van Dijk, about 30% of MNEs, accounting for roughly 90% of total profits, exceeded this threshold in 2018–2021 and were thus covered by the GMT. [^4] See Supplemental Appendix A. This figure aligns with the OECD’s estimate of 90% profit coverage (OECD, , 2020, # 505, p. 233). Nevertheless, over 300 billion EUR of MNEs’ profits remain outside its scope. Moreover, Orbis tends to oversample large firms (Bajgar et al., , 2020), so the 90% coverage rate should be viewed as an upper bound.
Does this incomplete coverage affect the effectiveness of the GMT? It does not if low-tax countries apply the GMT rate to all firms regardless of size. It does matter, however, if they differentiate their tax rates, applying the GMT rate to firms above the threshold and a lower rate to those below. Such a split is inherent in the regulation: tax havens may implement the GMT by raising their (low) base tax rates only for MNEs above the threshold through a top-up tax. [^5] This is known as the Qualified Domestic Minimum Top-Up Tax (QDMTT) (Devereux, , 2023). This system has been adopted in Bahrain, Ireland, Switzerland and the UAE, among other countries with base statutory corporate tax rates below 15% (PwC, , 2025).
@@
Against this background, we ask when tax competition leads countries to differentiate tax rates for large and small MNEs under the GMT. Does a 15% GMT rate induce such a split, and how large are the resulting revenue and welfare gains? How do these outcomes change with broader GMT coverage?
@@
To answer these questions, we build a simple model where heterogeneous MNEs headquartered in a non-haven country shift profits to tax-haven countries. Countries maximise welfare—a weighted sum of tax revenues and, for the non-haven, the post-tax income of MNEs’ owners—and compete for the MNEs’ profits. In the first stage, both the non-haven and the havens decide whether to commit to a uniform tax rates for all MNEs. In the second stage, countries set tax rates, either uniform or split, with large MNEs taxed at least at the GMT rate. The final stage involves MNEs’ profit-shifting decisions.
@@
Our analysis shows that introducing the GMT with partial coverage generates a sequence of non-cooperative tax equilibrium regimes between the non-haven and the havens. When the GMT rate is binding but low, the havens commit to it for all MNEs to raise the non-haven’s uniform tax rate. As the GMT rate rises, the havens begin to split their rates, applying the GMT to large MNEs and a lower rate to small MNEs. For sufficiently high GMT rates, the non-haven also splits its rate, as limited competition for large MNEs’ profits shifts the focus towards small MNEs.
@@
We calibrate our model using estimates of profit shifting and examine the quantitative effects of GMT reforms. A 15% GMT rate induces the havens to split their tax rates, which is consistent with observed responses in some tax havens including Ireland and Switzerland. It benefits the non-haven while harms the havens, generating net global revenue (welfare) gains of 136 billion USD (83 bUSD), comparable to existing estimates (Baraké et al., , 2022; Hugger et al., , 2024). If the GMT covers all firms, the 15% rate does not produce a split tax system but increases revenues and welfare in the non-haven and globally even more. This global revenue gain would have to be set against the additional compliance costs for multinationals, however.
@@
We contribute to the recent theoretical studies on tax competition under a GMT. They ask how introducing a GMT affects national and global revenues and welfare in the presence of country heterogeneity (Johannesen, , 2022; Hebous and Keen, , 2023); real economic activities (Janeba and Schjelderup, , 2023; Schjelderup and Stähler, , 2024); and endogenous tax-enforcement levels (Hindriks and Nishimura, , 2025). All these models assume the GMT applies uniformly, ignoring the firm-size threshold below which firms are exempt from the GMT. Our model highlights this threshold, allowing countries to tax firms above and below it differently. [^6] This links our model to the literature on discriminatory tax competition, where countries can tax firms differently by their mobility (Janeba and Peters, , 1999; Keen, , 2001; Bucovetsky and Haufler, , 2007; Gaigné and Wooton, , 2011).
Our simple quantitative exercise complements recent evaluations of corporate taxation using trade and multinational models (Wang, , 2020; Ferrari et al., , 2023; Shen, , 2024; Dyrda et al., , 2024), where taxes are assumed to be exogenous and/or uniform across firms. By contrast, we abstract from firms’ real activities and focus on the strategic role of size-dependent taxes.
Two empirical papers (Bachas et al., , 2025; Clifford et al., , 2025) closely related to ours examine which firms fall within the scope of the GMT using firm/affiliate level data for 16 countries and Germany, respectively. Their findings that GMT revenue gains depend strongly on its coverage emphasizes the relevance of our analysis.
@@ 1.1 Setup
We consider tax competition between a non-haven country, indexed by $n$, and a finite set of $H$ symmetric tax-haven countries, indexed by $h$. There is a continuum of heterogeneous MNEs which are headquartered in the non-haven and have affiliates in all havens. Headquarters earn (‘true’) exogenous profits $\pi\in[\underline{\pi},\infty)$ with distribution $F(\pi)$, arising from real activity in the non-haven. [^7] Firm heterogeneity could result from (executives’) productivity (e.g., Melitz, , 2003; Lippi and Schivardi, , 2014). MNEs can shift some of these profits to the havens to maximise their post-tax profits. In the following, we assume collective decision-making by the havens. This assumption is relaxed in the quantitative analysis later (see also footnote 16) and further examined in Supplemental Appendix B.
We assume the non-haven and all havens have signed the GMT agreement and must set tax rates no lower than the GMT rate $t_{M}$. [^8] Devereux, (2023) analyses how the GMT’s institutional design, especially the Under-Taxed Payments Rule (UTPR), incentivizes non-haven and haven countries to join once a critical mass of large non-havens participates. The GMT applies only to MNEs with profits $\pi\geq\pi_{M}$ (‘large MNEs’), allowing countries to set lower rates for firms with $\pi<\pi_{M}$ (‘small MNEs’).
@@
We analyse a three-stage game. In the first stage, the non-haven and havens simultaneously decide whether to commit to a single tax rate for all MNEs—at or above the GMT rate—or to allow differentiated rates. [^9] The ability to commit, a realistic feature of international tax relations, is standard both in the studies cited in footnote 6 and in related work on endogenous timing in games (Huck et al., , 2002; Kempf and Rota-Graziosi, , 2010). The choice between a uniform and a split tax system is more fundamental, and less frequently revised, than the choice of tax rates. The havens decide collectively on whether to unify or split their rates. In the second stage, countries set their tax rates non-cooperatively. Those that committed choose a single rate; others set different rates for large and small MNEs. In the final stage, MNEs shift profits given these tax rates. The game is solved by backward induction, yielding a sub-game perfect Nash equilibrium, referred to as the tax-competition equilibrium.
@@ 1.2 Profit shifting by multinationals
Each MNE shifts a share $\theta_{h}$ of its pre-tax profits $\pi$ from its non-haven headquarters to each tax haven affiliate $h$, and incurs transaction costs $C_{h}(\cdot)$ for doing so. These costs can be interpreted as concealment costs or as expected fines, if detected. We assume the costs are quadratic in the level of profit shifting but decrease with the size of a firm: $C_{h}(\theta_{h}\pi)=\delta(\theta_{h}\pi)^{2}/(2\pi)=\delta\theta_{h}^{2}\pi/2$, resulting in total shifting costs of $\sum_{h=1}^{H}C_{h}(\theta_{h}\pi)$. [^10] This specification is standard in the literature (e.g., Hines and Rice, , 1994; Huizinga and Laeven, , 2008; Suárez Serrato, , 2018; Bilicka et al., , 2024). The parameter $\delta$ captures the difficulty of profit shifting. [^11] This can be higher, for example, due to taxpayer audit (Lediga et al., , 2023).
Given the tax rates $t_{n}$ and $t_{h}$ respectively in the non-haven and each of the havens, an MNE maximises its global post-tax profits net of profit-shifting costs, denoted by $\widetilde{\pi}$, by choosing the share of profit shifting $\theta_{h}$ to each haven: [^12] The GMT tax base equals that of national corporate taxes, ignoring additional deductions (‘substance-based income exclusion’), an issue examined by Schjelderup and Stähler, (2024).
@@
The optimal $\theta_{h}$ in (2) implies that MNEs of different sizes share the same tax base elasticity. Substituting (2) into (1) and using the symmetry of havens, the tax base of an MNE with $\pi$ in the non-haven becomes $TB_{n}=[1-H(t_{n}-t_{h})/\delta]\pi$, yielding $\varepsilon_{n}=-\frac{dTB_{n}/TB_{n}}{dt_{n}/t_{n}}=\frac{t_{n}}{\delta-H(t_{n}-t_{h})}$, independent of $\pi$. Equal tax base elasticities remove countries’ incentives for tax discrimination across MNEs in the absence of the GMT. This allows us to focus on the splitting responses to a GMT that is applied only to large firms.
The empirical evidence on the relationship between the size of firms and their responsiveness to tax is inconclusive. Media coverage of tax avoidance by very large multinationals gives an impression that bigger firms have a higher tax-base elasticity. However, there are several studies which find the opposite result that the tax-base elasticity is indeed higher for small firms (not necessarily MNEs). [^13] Making use of bunching in the distribution of taxable income of firms in the UK, Devereux et al., (2014) find that small-sized firms change their taxable income more significantly in response to statutory tax rate changes, as compared to medium-sized firms. Applying a similar empirical strategy to US firms, Coles et al., (2022) find that the elasticity of taxable income with respect to effective tax rates is monotonically decreasing in firm size. See also Auliffe et al., (2023) and the references therein for the heterogeneous tax elasticities of tangible asset investments among firms of different sizes. Therefore, assuming the tax base elasticity to be independent of firm size may be considered a useful and not unrealistic benchmark.

### v7

@@
Abstract
@@
The Global Minimum Tax (GMT) is applied only to firms above a certain size threshold, permitting countries to set differential tax rates for small and large firms. We analyse tax competition among multiple tax havens and a non-haven country for heterogeneous multinationals to evaluate the effects of this partial coverage of GMT. Upon the introduction of a low but binding GMT rate, the havens commit to the single uniform GMT rate for all multinationals. However, gradual increases in the GMT rate induce the havens, and subsequently the non-haven, to adopt discriminatory, lower tax rates for small multinationals. Our calibration exercise shows that introducing a GMT rate close to 15% results in a regime where only the havens adopt split tax rates. Welfare and tax revenues fall in the havens but rise in the non-haven, yielding a positive net gain worldwide.
JEL classification: F23; F25; H87
Keywords: Multinational firms; Tax avoidance; Global minimum tax; Profit shifting
1 Introduction
@@
Profit shifting by multinational enterprises (MNEs) remains a central challenge for international corporate taxation. Aggregate estimates suggest that more than one-third of foreign-earned corporate profits are shifted to tax havens (Tørsløv et al., 2023). [^1] Combining micro- and macro-level data on portfolio investment, Coppola et al. (2021) highlight the role of tax havens as conduits through which firms access international capital markets. See Camarero et al. (2025) for an application of their dataset in the context of the EU. The important role of tax havens for profit shifting is confirmed in other studies using microdata of multinational affiliates in France (Davies et al., 2018), the U.K. (Bilicka, 2019), and the U.S. (Guvenen et al., 2022). In response to the large revenue losses caused by profit shifting, the OECD has launched an action plan to fight base erosion in OECD countries, and in particular the profit shifting to tax havens (OECD, 2013). A core development in this endeavor is the introduction of a global minimum tax (GMT), also known as Pillar Two. [^2] See Fuest and Neumeier (2023); Hebous et al. (2024) for the institutional background. Another outcome of the BEPS initiative is Pillar One, which reallocates taxing rights from the country of a MNE’s residence to the market jurisdiction (Mukunoki et al., 2025; Ogawa and Tsuchiya, 2025; Richter, 2025). Pillar One aims to address the tax challenges posed by the digital economy (see Lassmann et al., 2025 for how multinational digital platforms respond to corporate taxes). In 2021, more than 130 countries including many tax havens agreed on a 15% GMT rate; by August 2025, 65 countries had drafted or enacted legislation implementing the GMT. [^3] See Tax Foundation, ‘The Latest on the Global Tax Agreement’: https://taxfoundation.org/blog/global-tax-agreement/ (accessed January 29, 2026). The United States had already implemented, in its 2017 Tax Cuts and Jobs Act, a tax on Global Intangible Low-Tax Income (GILTI) that works in many respects like a GMT (Chodorow-Reich et al., 2024).
@@
An important limitation of the GMT is that it applies only to ‘large’ multinationals, defined as groups with at least 750 million EUR in annual revenues in two of the last four years. [^4] Since 2016, such groups have been required to file country-by-country reports (CbCR) on their activities, profits, and taxes in each country of operation (Doeleman et al., 2024, Section 2). The revenue threshold reflects the balance between limiting profit shifting and minimising compliance costs (OECD, 2022, p.14). Fig. 1 shows that, according to the Orbis database of Bureau van Dijk, about 30% of MNEs, accounting for roughly 90% of total profits, exceeded this threshold in 2018–2021 and were thus covered by the GMT. [^5] This figure aligns with the OECD’s estimate of 90% profit coverage (OECD, 2020, # 505, p. 233). Nevertheless, over 300 billion EUR of MNEs’ profits remain outside its scope. Moreover, Orbis tends to oversample large firms (Bajgar et al., 2020), and hence the 90% coverage rate should be seen as an upper bound.
One implication of this incomplete coverage is that low-tax countries might respond to the introduction of the GMT by using a split corporate tax system, where a lower tax rate applies to firms below the GMT threshold. In fact, such a split is inherent in the regulation: tax havens can implement the GMT by raising their (low) base tax rates only for MNEs above the threshold through a specific top-up tax. [^6] This is known as the Qualified Domestic Minimum Top-Up Tax (QDMTT). See Devereux (2023) for a discussion. Tax havens are fully aware of this option. Ireland, for example, has decided to keep its general tax rate at 12.5%, but top up the tax rate to 15% for affiliates of foreign MNEs above the threshold (Government of Ireland, 2023). Several other major low-tax countries, including Bahrain, Bulgaria, Hungary, Switzerland and the UAE have adopted similar split tax systems by joining the GMT agreement while maintaining their base statutory corporate tax rates below 15% (PwC, 2025). [^7] The latest corporate tax rates around the world can be found in ‘Tax Foundation, ‘Corporate Tax Rates Around the World, 2025’: https://taxfoundation.org/data/all/global/corporate-tax-rates-by-country-2025/ (accessed January 29, 2026). Note that not all low-tax countries have adopted a split tax system: Cyprus and Gibraltar, for example, raised their base rates from 12.5% to 15% for both large and small MNEs, effective in 2026 and 2024, respectively. See Shanda Consult, ‘The Cyprus Tax Reform 2026: A Strategic Evolution for Business and Compliance’: https://shandaconsult.com/cyprus-tax-reform-2026 and PwC, ‘Gibraltar—Taxes on corporate income’: https://taxsummaries.pwc.com/gibraltar/corporate/taxes-on-corporate-income (accessed January 29, 2026).
Fig. 1. GMT coverage.
Source: Orbis database, own calculations.
Notes: The left bar in each year shows the share of MNEs covered by the GMT out of all MNEs. The right bar in each year shows the share of pre-tax profits earned by MNEs covered by GMT out of pre-tax profits by all MNEs. The MNEs covered by GMT are those whose annual revenues exceed $750$ million EUR in at least two of the last four years out of all MNEs. See A for details.
@@
These responses show that the effects of the GMT cannot be analysed assuming a single, uniform corporate tax rate across all countries. Against this background, this paper asks when tax competition leads countries to differentiate tax rates for large and small MNEs under the GMT, whether a 15% GMT rate induces such a split, and how large are the resulting revenue and welfare gains.
@@
To answer these questions, we build a simple model where heterogeneous MNEs headquartered in a non-haven country shift profits to tax-haven countries. Countries maximise welfare—a weighted sum of tax revenues and, for the non-haven, the post-tax income of MNEs’ owners—and compete for the MNEs’ profits. In the first stage, both the non-haven and the havens decide whether to commit to a uniform tax rate for all MNEs. In the second stage, countries set tax rates, either uniform or split, with large MNEs taxed at least at the GMT rate. The final stage involves MNEs’ profit-shifting decisions.
@@
Our analysis shows that introducing the GMT with partial coverage generates a sequence of non-cooperative tax equilibrium regimes between the non-haven and the havens. When the GMT rate is binding but low, the havens commit to it for all MNEs to raise the non-haven’s uniform tax rate. As the GMT rate rises, the havens start splitting their rates, applying the GMT to large MNEs and a lower rate to small MNEs. For sufficiently high GMT rates, the non-haven also splits its rate, as limited competition for large MNEs’ profits shifts the focus towards small MNEs.
@@
We then calibrate our model using estimates of profit shifting and examine the quantitative effects of GMT reforms. For our basic model, a 15% GMT rate leads both the non-haven and the havens to choose uniform rates for all firms, whereas a 16% GMT rate induces the havens to split their tax rates. In an extended model where tax havens choose their tax rates in a decentralized way, a 15% GMT rate is sufficient to induce a split tax system for the havens in equilibrium. This splitting result is consistent with the observed responses in several tax havens, including Ireland and Switzerland, as discussed earlier. Our quantitative results also show that introducing a 15% GMT rate benefits the non-haven while harming the havens, generating net global revenue gains of 149–156 billion USD (and welfare gains of 92–96 billion USD). These revenue gains are comparable to existing estimates in the literature (Baraké et al., 2022; Hugger et al., 2024).
@@
Our analysis contributes to the recent theoretical studies on tax competition under a GMT. They ask how introducing a GMT affects national and global revenues and welfare in the presence of country heterogeneity (Johannesen, 2022; Hebous and Keen, 2023); real economic activities (Janeba and Schjelderup, 2023; Schjelderup and Stähler, 2024); and endogenous tax-enforcement levels (Hindriks and Nishimura, 2025). All these models assume the GMT applies uniformly, ignoring the firm-size threshold below which firms are exempt from the GMT. Our model highlights this threshold, allowing countries to tax firms above and below it differently. [^8] This links our model to the literature on discriminatory tax competition, where countries can tax firms differently by their mobility (Janeba and Peters, 1999; Keen, 2001; Peralta et al., 2006; Gaigné and Wooton, 2011; Haufler and Runkel, 2012).
Our simple quantitative exercise yields results that are comparable to more detailed estimates of the revenue effects of a GMT (Baraké et al., 2022; Hugger et al., 2024). It also complements recent evaluations of corporate taxation using trade and multinational models (Wang, 2020; Ferrari et al., 2023; Shen, 2024; Dyrda et al., 2024; Santacreu, 2026), where taxes are assumed to be exogenous and/or uniform across firms.By contrast, we abstract from firms’ real activities and focus on the strategic role of size-dependent taxes. One of the few papers offering empirical evidence on minimum tax regimes is Buettner and Poehnlein (2024), who study the effects of introducing a minimum local business tax rate in Germany. Finally, two recent empirical papers (Bachas et al., 2025; Clifford et al., 2025) examine which firms fall within the scope of the GMT using firm/affiliate level data for 16 countries and Germany, respectively. Their findings that GMT revenue gains depend strongly on its coverage emphasize the relevance of our analysis.
We proceed as follows. In Section 2 we set up our model of tax competition. Section 3 characterises the sub-game perfect Nash equilibria that result in different tax regimes. Section 4 turns to the welfare effects of introducing a GMT, and of increasing its tax rate and coverage. Section 5 calibrates our model to real-world data and quantifies its effects. Section 6 concludes.
@@ 2.1 Setup
We consider tax competition between a non-haven country, indexed by $n$, and a finite set of $H$ symmetric tax-haven countries, indexed by $h$. There is a continuum of heterogeneous MNEs which are headquartered in the non-haven and have affiliates in all havens. [^9] To focus on the aggregate level of profit shifting relevant for countries’ corporate tax revenues, we abstract from the endogenous location choice of MNEs. On the evolution of US-owned MNEs over time and space, see Guvenen et al. (2022) for profit shifting and Garetto et al. (2024) for real economic activities. Headquarters earn (‘true’) exogenous profits $\pi\in[\underline{\pi},\infty)$ with distribution $F(\pi)$, arising from real activity in the non-haven.MNEs can shift some of these profits to the havens to maximise their post-tax profits. In the following, we assume (i) true profits $\pi$ independent of taxes and (ii) the collective decision-making by the havens for analytical simplicity. These assumptions are relaxed in the quantitative analysis in Section 5.2, where we explore the robustness of our results.
We assume the non-haven and all havens have signed the GMT agreement and must set tax rates no lower than the GMT rate $t_{M}$. [^10] Devereux (2023) analyses how the GMT’s institutional design, especially the Under-Taxed Payments Rule (UTPR), incentivizes non-haven and haven countries to join once a critical mass of large non-havens participates. The GMT applies only to MNEs with profits $\pi\geq\pi_{M}$ (‘large MNEs’), allowing countries to set lower rates for firms with $\pi<\pi_{M}$ (‘small MNEs’).
@@
We analyse a three-stage game. In the first stage, the non-haven and havens simultaneously decide whether to commit to a single tax rate for all MNEs that is at or above the GMT rate, or to allow differentiated rates. [^11] The ability to commit, a realistic feature of international tax relations, is standard both in the studies cited in footnote 8 and in related work on endogenous timing in games (Huck et al., 2002; Kempf and Rota-Graziosi, 2010). The choice between a uniform and a split tax system is more fundamental, and less frequently revised, than the choice of tax rates. The havens decide collectively on whether to unify or split their rates. In the second stage, countries set their tax rates non-cooperatively. Those that committed choose a single rate; others set different rates for large and small MNEs. In the final stage, MNEs shift profits given these tax rates. The game is solved by backward induction, yielding a sub-game perfect Nash equilibrium. We also refer to this as the tax-competition equilibrium.
@@ 2.2 Profit shifting by multinationals
Each MNE shifts a share $\theta_{h}$ of its pre-tax profits $\pi$ from its non-haven headquarters to each tax haven affiliate $h$, and incurs transaction costs $C_{h}(\cdot)$ for doing so. These costs can be interpreted as concealment costs or as expected fines, if detected. We assume the costs are quadratic in the level of profit shifting but decrease with the size of a firm: $C_{h}(\theta_{h}\pi)=\delta(\theta_{h}\pi)^{2}/(2\pi)=\delta\theta_{h}^{2}\pi/2$, resulting in total shifting costs of $\sum_{h=1}^{H}C_{h}(\theta_{h}\pi)$. [^12] This specification is standard in the literature (e.g., Hines and Rice, 1994; Huizinga and Laeven, 2008; Suárez Serrato, 2018; Bilicka et al., 2024). The parameter $\delta$ captures the difficulty of profit shifting.
Given the tax rates $t_{n}$ and $t_{h}$ respectively in the non-haven and each of the havens, an MNE maximises its global post-tax profits net of profit-shifting costs, denoted by $\widetilde{\pi}$, by choosing the share of profit shifting $\theta_{h}$ to each haven: [^13] The GMT tax base equals that of national corporate taxes. We thus ignore additional deductions (‘substance-based income exclusion’), an issue examined by Schjelderup and Stähler (2024).
@@
The optimal $\theta_{h}$ in (2) implies that MNEs of different sizes share the same tax base elasticity. Substituting (2) into (1) and using the symmetry of havens, the tax base of an MNE with $\pi$ in the non-haven becomes $TB_{n}=[1-H(t_{n}-t_{h})/\delta]\pi$, yielding $\varepsilon_{n}=-\frac{dTB_{n}/TB_{n}}{dt_{n}/t_{n}}=\frac{t_{n}}{\delta-H(t_{n}-t_{h})}$, which is independent of $\pi$. Equal tax base elasticities thus remove countries’ incentives to tax-discriminate between MNEs in the absence of the GMT. This allows us to focus on the splitting responses to a GMT that is applied only to large firms.
The empirical evidence on the relationship between the size of firms and their responsiveness to tax is not conclusive. Media coverage of tax avoidance by very large multinationals gives an impression that bigger firms have a higher tax-base elasticity. However, there are several studies which find the opposite result that the tax-base elasticity is indeed higher for small firms (not necessarily MNEs). [^14] Making use of bunching in the distribution of taxable income of firms in the UK, Devereux et al. (2014) find that small-sized firms change their taxable income more significantly in response to statutory tax rate changes, as compared to medium-sized firms. Applying a similar empirical strategy to US firms, Coles et al. (2022) find that the elasticity of taxable income with respect to effective tax rates is monotonically decreasing in firm size. See also Auliffe et al. (2023) and the references therein for the heterogeneous tax elasticities of tangible asset investments among firms of different sizes. A detailed analysis of French MNEs using Country-by-Country Reports (CbCR) empirically supports our result in eq. (2) that firms of different size shift roughly the same share of their profits (Aliprandi et al., 2025, Figure 5). This is compatible with the finding that large MNEs are responsible for most of the total amount of profit shifting (Clifford et al., 2025), given their large share of total profits. In summary, assuming the tax base elasticity to be independent of firm size seems to be a useful and empirically plausible baseline.